OpenStudy (anonymous):

The perimeter of an rectange is 380, the Length is 90 yd longer than the width, what are the dimensions

Directrix (directrix):

Let W be the width of the rectangle. Then, W + 90 is the length of the rectangle.

Directrix (directrix):

A rectangle has 4 sides. Opposite sides are congruent and have the same measure. Perimeter is the sum of the distances around the rectangle.

OpenStudy (anonymous):

w+90=380 w=290?

Directrix (directrix):

So, W + W + W + 90 + W + 90 = 380

Directrix (directrix):

> w+90=380 w=290? No.

OpenStudy (anonymous):

3w+90=390

OpenStudy (anonymous):

3w=300

OpenStudy (anonymous):

w=100

Directrix (directrix):

W + W + W + 90 + W + 90 = 380 4W + 180 = 380

OpenStudy (anonymous):

okay according to your problem its 50

Directrix (directrix):

I think your equation was not correct.

OpenStudy (anonymous):

the steps i added the w's, then subtracted and divide

Directrix (directrix):

If the width is 50, then what is the length?

OpenStudy (anonymous):

140

OpenStudy (anonymous):

because it says 90 longer

Directrix (directrix):

Correct.

Directrix (directrix):

You can check the answers: 50 + 50 + 140 + 140 = 380.